# Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate

### v1

@@ 5 Summary and Future Directions
We have announced the detection of the first free-floating planet candidate found with TESS data. Identified in a search of Sector 61 FFI light curves for short-duration microlensing events, the event is associated with TIC-107150013, a source sitting at $d_{s}=3.193$ kpc with radius $R_{s}=12.9\,R_{\odot}$. Fitting the observed signal with a four-parameter finite-source lensing model yields an Einstein-crossing time of $t_{E}=0.074^{+0.002}_{-0.002}$ days and finite-source parameter $\rho=4.55^{+0.08}_{-0.07}$, implying an Einstein radius of $\theta_{E}=4.1^{+0.4}_{-0.5}\,\mu\text{as}$. This value places the lens firmly in the terrestrial-mass range, with a mass less than $10~{}M_{\oplus}$ if within $\sim 2.6$ kpc, and less than $1~{}M_{\oplus}$ if within $\sim 1$ kpc.
@@
This initial result demonstrates that TESS has the opportunity to significantly improve our understanding of FFPs with terrestrial and sub-terrestrial masses, a range of the FFP mass spectrum that is poorly-constrained and will remain challenging to observe even in future surveys, such as the Roman Space Telescope’s Galactic Time Domain Survey (GBTDS) (Gaudi, 2022). Furthermore, the search presented here is preliminary, having been based on a small and relatively bright subset of all TESS target stars. We are currently repeating our search across all TESS sectors, and plan to use a fainter dataset of TGLC light curves down to $T=16$ mag. This combination will increase the number of possible source stars by over two orders of magnitude. These data will allow us to confirm the existence of the population suggested by the observed FFP candidate in Sector 61, as well as to characterize its origin. The results of these ongoing searches will help inform theoretical studies of formation mechanisms for rogue planets, and will play a crucial role in calibrating the expected yield and search strategy of the GBTDS. Though never intended as a microlensing satellite, TESS is opening a window to the dark, enigmatic population of worlds that lie drifting between the stars.
Acknowledgements
We thank Te Han for discussions about TGLC light curves, and Avi Shporer for discussions about follow-up prospects for our evolving candidates. MK acknowledges support by the Juan Carlos Torres postdoctoral fellowship from the MIT Kavli Institute for Astrophysics and Space Research. WD and NS acknowledge the support of DOE grant No. DE-SC0010107.
@@ Acknowledgements
This paper includes data collected by the TESS mission, which are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Data Availability
All 1.3 million QLP light curves from Sector 61 (and additional QLP light curves used for this work, including the 6.2 million light curves from Sectors 62 – 65) are publicly available as High-Level Science Products (HLSPs) on MAST at https://archive.stsci.edu/hlsp/qlp. Our additional TGLC light curve for TIC-107150013 was extracted using TGLC version 0.6.0, following the tutorial detailed at https://github.com/TeHanHunter/TESS_Gaia_Light_Curve. We set the target ID to TIC 107150013, FFI cut-out size to 50, and sector number to 61.

### v2

@@ 5 Summary and Future Directions
The search presented here is preliminary, having been based on a small and relatively bright subset of all TESS target stars. We are currently repeating our search across all TESS sectors, and plan to use a fainter dataset of TGLC light curves down to $T=16$ mag. This combination will increase the number of possible source stars by over two orders of magnitude. These data will provide the opportunity to continue searching for FFP signals in the sub-terrestrial mass range. Furthermore, our ongoing search for such signals will provide key insight on possible sources of false positive contamination for future space-based missions such as the Roman Space Telescope’s Galactic Bulge Time Domain Survey (GBTDS) (Akeson et al., 2019; Penny et al., 2019) and Earth 2.0 (Ge et al., 2022).
Our initial results presented here demonstrate that TESS can be a powerful tool to search for signals like that of free-floating planets at masses below that of Earth. The results of our ongoing search will play a key role in calibrating the expected yield and search strategy of the GBTDS. Though never intended as a microlensing satellite, TESS is helping to open a window to the dark, enigmatic population of worlds that lie drifting between the stars.
@@
We thank the referee for their useful feedback which significantly improved the manuscript. We thank J.P. Beaulieu for his useful input on revising this manuscript, Marc Hon for discussions about giant stars and activity, Te Han for discussions about TGLC light curves, Avi Shporer for discussions about follow-up prospects for our events, and Jon Jenkins for discussions about SPOC 2-min and 20-sec cadence light curves. MK acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2024-06452. Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), RGPIN-2024-06452. MK also acknowledges support by the Juan Carlos Torres postdoctoral fellowship from the MIT Kavli Institute for Astrophysics and Space Research. WD was supported by NSF grant PHY-2210361 and the Maryland Center for Fundamental Physics. WD and NS also acknowledge the support of DOE grant No. DE-SC0010107. BSG was supported by NASA Grant 80NSSC24M0022 and by the Thomas Jefferson Chair Endowment for Discovery and Space Exploration. This paper includes data collected by the TESS mission, which are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.